# Battery Capacity Knee-Onset Identification and Early Prediction Using Degradation Curvature

### v2

@@ 4.1.3 Synthetic dataset
Figure 6: The knee-onset and knee identification results of the cylindrical LGM50 cell.
The loss of capacity is solely caused by LLI due to SEI growth on the normal particle surface and on the cracked surfaces. For a large cracking rate (i.e., 30 times the standard particle cracking rate purewal2014degradation ), the LLI begins with a square root dependence on time until it reaches the inflection point, which is close to the identified capacity knee-onset point, as shown in Fig. 7. After the inflection point, the LLI accelerates exponentially as the cracks propagate. The accelerated LLI as the internal state change is the only cause for the simulated knee occurrence in this case.
@@
The inconsistent results of using the double Bacon-Watts model to identify knee-onset and knee on two experimental battery degradation datasets indicate a lack of generalizability of the double Bacon-Watts model towards various battery chemistries for a wide range of operating conditions. In contrast, the generalizability of our proposed capacity knee identification method has been demonstrated on the two experimental degradation data of two battery chemistry types under a wide range of operating conditions.
4.2 Results of knee-onset early prediction
As shown in Fig. 3 and 4, both knee and knee-onset points identified using our proposed method show strong correlations with the end of life. Additionally, compared to the knee alone, the knee-onset can give a much earlier warning of accelerated degradation (note that on average there are 323 cycles for the LFP cells and 280 cycles for the NMC cells between the knee-onset and the identified knee). Each charge-discharge cycle takes 50 min on average, and consequently, a reduction of 323 cycles in the Toyota Research Institute dataset translates to a reduction of experimental time by approximately 269 hours. Therefore, learning an early-prediction model for battery knee-onset prediction requires much less degradation data than that for battery knee prediction, which in turn significantly reduces experimental time and its associated costs.
Most of knee-onset points identified using the proposed method fall between 150 and 270 cycles in the Toyota Research Institute dataset. Thus, these 169 cells are first graded into three classes, i.e., early-knee-onset class ($<150\ \text{cycles}$), normal-knee-onset class ($150-270\ \text{cycles}$), and late-knee-onset class ($>270\ \text{cycles}$). Then, in order to learn a generalized knee-onset early-prediction model, the stratified random sampling method reitermanova2010data is employed to randomly split 169 cells, with 80% in a training set and 20% in a test set. Equal ratios of early-knee-onset cells, normal-knee-onset cells, and late-knee-onset cells are preserved in the training and test set at each split. Furthermore, the stratified random sampling is repeated 5 times in order to reduce the random effect of the train-test split. The final model performance is averaged over 5 train-test splits.
For any machine learning model, input features that are extracted from early degradation data determine their prediction performance. It has been demonstrated by Severson et al. severson2019data that a 6-feature set (see Table. 5) results in an early-prediction model with the best battery lifetime prediction performance among all 3 feature sets (i.e., 1-feature set, 6-feature set, and 9-feature set). Therefore, this 6-feature set is used here for training knee-onset early-prediction models and predicting knee-onset points later on. We then select the gradient boosting regression tree (GBRT) model to map input features to knee-onset points, thanks to its advantages, such as capturing nonlinear relationships, robustness to outliers, and high predictive accuracy.
@@ 4.2 Results of knee-onset early prediction
A sensitivity analysis of the model performance using different amounts of early degradation data (from the first 15 cycles to the first 35 cycles) is illustrated in Fig. 8. It can be seen that the prediction errors, measured by root-mean-square error (RMSE) and mean absolute percentage error (MAPE), significantly decrease with increasing number of cycles. The lowest prediction errors are obtained using degradation data from the first 30 cycles, after which the prediction errors do not decrease significantly. Thus, this model is capable of providing knee-onset prediction with high accuracy after only 30 cycles. Moreover, compared to battery lifetime early prediction using the first 100 cycles data in Severson et al. severson2019data , a reduction of 70 cycles in battery knee-onset early prediction translates to a reduction of experimental time by approximately 58 hours, which in turn reduces experimental costs.
@@ 5 Conclusions
Throughout the literature review, knees indicating the beginning of accelerated degradation and possible safety issues have been observed to occur within a window of 70-95 % of the initial nominal capacity in experimental testing of commercial lithium-ion cells for various chemistries and different operating conditions, which To prepare for a successful second-life battery market, concerns arising from possible knee occurrence during first-life and second-life need to be addressed at the repurposing stage.
@@
As the first step to address the aforementioned concerns, the root causes for the formation of battery capacity knees have been considered and a curvature-based method to identify capacity knee and capacity knee-onset on the capacity fade curve is proposed. Specifically, the approximated curvature was first used to represent the rate of change of capacity fade in discrete time, from which a new degradation phenomenon was found. The capacity knee-onset and capacity knee were identified as the start and the end points, respectively, of a second state of the degradation process. It was found that the approximated curvature fluctuated significantly in this state. The proposed capacity knee identification method was benchmarked to the state-of-the-art knee identification method (i.e., the double Bacon-Watts model) on both synthetic degradation data and experimental degradation data of both LFP and NMC cells. In the results for the NMC cells, it was demonstrated that the state-of-the-art method failed to identify the knee on the capacity fade curve while our proposed capacity knee identification method successfully identified the knee. The results of the capacity knee identification on synthetic degradation data further validated the effectiveness of our proposed capacity knee identification method. In contrast to capacity knee identification alone, the knee-onset can give a much earlier warning of accelerated degradation (an average of 323 cycles was found for the LFP cells and an average of 280 cycles for the NMC cells between the knee-onset and the knee). Therefore, learning early-prediction models for battery knee-onset prediction can therefore significantly reduce experimental time and its associated costs. Furthermore, knee-onset prediction can have significant economic value in various industrial applications, such as battery grading, battery replacement planning, and battery repurposing to second-life applications. The capacity knee-onsets and capacity knees that are identified can also be used to systematically evaluate the knee-related prediction performance of both model-based methods and data-driven methods.
@@
Our proposed capacity knee identification method has been validated on battery degradation data from static cycling tests. Battery degradation data from dynamic cycling tests under realistic driving profiles are also suggested to be investigated. In an electric vehicle, we may not have access to all the cells, but at least the cell with the lowest capacity is accessible, and therefore it would be recommended to validate our proposed capacity knee identification method on capacity fade data that is real-time estimated in the field in the next step. Lastly, considering the close relationship between capacity knees and resistance/impedance elbows, it would also be interesting to adapt the proposed capacity knee identification method accordingly so that the internal resistance elbows can be identified and predicted from internal resistance/impedance data.

### v3

@@ 4.1.3 Validation on synthetic dataset
Figure 6: The knee-onset and knee identification results of the cylindrical LGM50 cell 2.
The loss of capacity is solely caused by LLI due to SEI growth on the normal particle surface and on the cracked surfaces. For a large cracking rate (i.e., 30 times the standard particle cracking rate [39]), the LLI begins with a square root dependence on time until it reaches the inflection point, which is close to the identified capacity knee-onset point, as shown in Fig. 7. After the inflection point, the LLI accelerates exponentially as the cracks propagate. The accelerated LLI as the internal state change is the only cause for the simulated knee occurrence in this case. The use of synthetic data has the potential for further analysis into the interactions between degradation mechanisms and the evolution of degradation modes behind the observed knee phenomena. Other degradation pathways that may consist of state trajectories of multiple degradation modes can also lead to a knee on the capacity fade curve - knee pathways [6]. Therefore, to enable battery degradation diagnosis including knee identification, a larger synthetic dataset that covers a range of other knee pathways needs to be generated, or found in the literature, for example, the Hawaii Natural Energy Institute (HNEI) synthetic dataset by Dubarry and Beck [40] [41].
@@
Overall, the inconsistent results of using the state-of-the-art model (i.e., double Bacon-Watts model) to identify knee-onset and knee on two experimental battery degradation datasets and one synthetic dataset indicate a lack of generalizability of the double Bacon-Watts model towards various battery chemistries for a wide range of operating conditions. In contrast, the generalizability of our proposed capacity knee identification method has been demonstrated on three battery chemistry types, two experimental degradation datasets, and one synthetic dataset, under a wide range of operating conditions. Moreover, as knee identification is formulated as an unsupervised learning problem, knee labeling is not required beforehand. Exact knee-onset and knee points on the capacity fade curve are identified using capacity data as the only input, which is preferable in practical applications, such as a systematic evaluation of the knee prediction performance of both model-based methods and data-driven methods, and facilitating classification of retired electric vehicle batteries from safety and performance perspectives.
4.2 An application case study: knee-onset early prediction
As shown in Fig. 3 and 4, both knee and knee-onset points identified using our proposed method show strong correlations with the end of life. Additionally, compared to the knee alone, the knee-onset can give a much earlier warning of accelerated degradation (note that on average there are 323 cycles for the LFP cells and 280 cycles for the NMC cells between the knee-onset and the identified knee). Each charge-discharge cycle takes 50 min on average, and consequently, a reduction of 323 cycles in the TRI dataset translates to a reduction of experimental time by approximately 269 hours. Therefore, learning an early-prediction model for battery knee-onset prediction requires much less degradation data than that for battery knee prediction, which in turn significantly reduces experimental time and its associated costs.
4.2.1 Feature engineering
To develop machine learning models for battery knee-onset early prediction, input features that are extracted from early degradation data determine their prediction performance. It has been demonstrated by Severson et al. [23] that a 6-feature set (see Table. 5) results in an early-prediction model with the best battery lifetime prediction performance among the 3 feature sets investigated (i.e., 1-feature set, 6-feature set, and 9-feature set). Therefore, this 6-feature set is used here for training knee-onset early-prediction models and predicting knee-onset points later on.
@@ 4.2.1 Feature engineering
4.2.2 Train-test split
Most of knee-onset points identified using the proposed method fall between 150 and 270 cycles in the TRI dataset. Thus, these 169 cells are first graded into three classes, i.e., early-knee-onset class ($<150\ \text{cycles}$), normal-knee-onset class ($150-270\ \text{cycles}$), and late-knee-onset class ($>270\ \text{cycles}$). Then, in order to learn a generalized knee-onset early-prediction model, the stratified random sampling method [42] is employed to randomly split 169 cells, with 80% in a training set and 20% in a test set. Equal ratios of early-knee-onset cells, normal-knee-onset cells, and late-knee-onset cells are preserved in the training and test set at each split. Furthermore, the stratified random sampling is repeated 5 times in order to reduce the random effect of the train-test split. The final model performance is averaged over 5 train-test splits.
4.2.3 Model selection
By combining a statistical technique called boosting, the Gradient boosting regression trees (GBRT) aggregate a set of ”weak” trees to form a single ”strong” tree. During the training stage, new trees are generated sequentially to correct the prediction errors of previous trees. This is achieved by minimizing a predefined loss function (e.g., least squares), which quantifies the difference between predicted and actual target values. Moreover, the contribution of each tree to the ensemble model is weighted by the learning rate to prevent overfitting [43]. It has been demonstrated that the GBRT model provided the best battery lifetime early prediction performance among other models (i.e., elastic net, support vector regression, random forests, gaussian process regression, quantile regression forests, quantile regression gradient boosting) using this 6-feature set in previous work [44]. Since the knee-onset points identified using our proposed method have shown strong correlations with the end of life (see Fig. 3), the GBRT model is selected for knee-onset early prediction in this case study.
4.2.4 Model performance evaluation
Firstly, a sensitivity analysis of the model performance using different amounts of early degradation data (from the first 15 cycles to the first 35 cycles) is illustrated in Fig. 8. It can be seen that the prediction errors, measured by root-mean-square error (RMSE) and mean absolute percentage error (MAPE), significantly decrease with increasing number of cycles. The lowest prediction errors are obtained using degradation data from the first 30 cycles, i.e., RMSE of 59.2 cycles and MAPE of 20.2%, after which the prediction errors do not decrease significantly. Thus, this model is capable of providing knee-onset prediction with high accuracy after only 30 cycles.
@@ 4.2.4 Model performance evaluation
To further investigate the outcomes of early knee-onset prediction using the proposed identification method and its correlation with the battery’s end of life, we illustrate the relationship between predicted knee-onset (using the first 30 cycles data) and identified knee-onset, predicted knee-onset and end of life in Fig. 9. Once more there are strong linear correlations between predicted knee-onset and knee-onset identified using our proposed method ($\rho=0.816$), and between predicted knee-onset and end of life ($\rho=0.816$).
Figure 9: The relationship between predicted knee-onset (using the first 30 cycles data) and identified knee-onset (left), predicted knee-onset and end of life (right) for 169 cells in the Toyota Research Institute dataset.
Moreover, compared to battery lifetime early prediction using the first 100 cycles data in Severson et al. [23], a reduction of 70 cycles in battery knee-onset early prediction translates to a reduction of experimental time by approximately 58 hours, which in turn reduces experimental costs.
@@ 5 Conclusions
Throughout our literature review, for various chemistries and different operating conditions, we find knees indicating the beginning of accelerated degradation and possible safety issues to occur within a window of 70-95 % of the initial nominal capacity in experimental testing of commercial lithium-ion cells. To prepare for a successful second-life battery market, concerns arising from possible knee occurrence during first-life and second-life need to be addressed at the repurposing stage.
@@
As the first step to address such concerns, the root causes for the formation of battery capacity knees have been considered and a curvature-based method to identify capacity knee and capacity knee-onset from the capacity fade curve is proposed. By analysis of the approximated curvature in discrete time, a new degradation phenomenon was found. The capacity knee-onset and capacity knee were identified as the start and the end points, respectively, of a transition state of the degradation process, where the approximated curvature fluctuates significantly. The proposed capacity knee identification method was benchmarked to the state-of-the-art knee identification method (i.e., the double Bacon-Watts model) on both synthetic degradation data and experimental degradation data of both LFP and NMC cells. In the results for the NMC cells, it was demonstrated that the state-of-the-art method failed to identify the knee on the capacity fade curve while our proposed capacity knee identification method successfully identified the knee. The results of the capacity knee identification on synthetic degradation data further validated the effectiveness of our proposed capacity knee identification method. In contrast to capacity knee identification alone, the knee-onset can give a much earlier warning of accelerated degradation (an average of 323 cycles was found for the LFP cells and an average of 280 cycles for the NMC cells between the knee-onset and the knee). Therefore, learning early-prediction models for battery knee-onset prediction can therefore significantly reduce experimental time and its associated costs. Furthermore, knee-onset prediction can have significant economic value in various industrial applications, such as battery grading, battery replacement planning, and battery repurposing to second-life applications. The capacity knee-onsets and capacity knees that are identified can also be used to systematically evaluate the knee-related prediction performance of both model-based methods and data-driven methods.
@@
Our proposed capacity knee identification method has been validated on battery degradation data from static cycling tests. For wider applications, the method needs to be evaluated on dynamic cycling test data, such as realistic driving profiles for electric vehicles. Then we may not have access to all the cells, but at least the cell with the lowest capacity should be accessible. Therefore it would be recommended to validate our proposed capacity knee identification method on capacity fade data that is real-time estimated in the field in a next step. The synthetic dataset of NMC cells with cracking-induced knees has been generated for validation purposes. The use of this synthetic dataset has demonstrated the potential of further analysis into the interactions between degradation mechanisms and the evolution of degradation modes that are associated with knee occurrence. Besides, the knee occurrence can also depend on other degradation pathways that may consist of state trajectories of multiple degradation modes. Hence, to enable online battery degradation diagnosis including knee detection, a larger synthetic dataset that covers a wide range of different knee pathways is needed. Lastly, considering the close relationship between capacity knees and resistance/impedance elbows, it would also be interesting to adapt the proposed capacity knee identification method accordingly so that the internal resistance elbows can be identified and predicted from internal resistance/impedance data.